# InSpatio-WorldFM: An Open-Source Real-Time Generative Frame Model

### 12 Mar 2026 v1

@@
Xiaoyu Zhang, Weihong Pan, Zhichao Ye, Jialin Liu, Yipeng Chen, Nan Wang,
Xiaojun Xiang, Weijian Xie, Yifu Wang, Haoyu Ji, Siji Pan, Zhewen Le, Jing Guo,
Xianbin Liu, Donghui Shen, Ziqiang Zhao, Haomin Liu, Guofeng Zhang
InSpatio Team
@@ 3 Evaluation
To enable real-time interaction, we apply distribution matching distillation to obtain a distilled model for low-latency deployment. In its baseline configuration, InSpatio-WorldFM runs at approximately 10 FPS at 512×512 resolution on a single NVIDIA A100 GPU with around 50-70 ms interaction latency. Owing to its low GPU memory consumption, it also enables interactive exploration on consumer-grade GPUs and achieves 7 fps on RTX 4090 with single-step inference. With engineering optimizations such as KV-cache management and efficient VAE latent caching, the inference speed can be further significantly improved. While distillation introduces a quality-speed trade-off, the perceptual difference remains minimal in practice. Visual comparison between the fundamental model and the distilled variant reveals that InSpatio-WorldFM successfully retains spatial coherence and geometric accuracy. The accelerated model continues to generate plausible content across diverse camera trajectories without noticeable degradation in structural consistency or the emergence of visual artifacts, confirming that our distillation approach effectively balances real-time performance with generation fidelity.
@@ 4.2 Future Work
5 Contributors
Model: Xiaoyu Zhang, Weihong Pan, Zhichao Ye, Yipeng Chen
Data: Jialin Liu, Xiaojun Xiang, Weijian Xie
Engineering: Nan Wang, Siji Pan, Jing Guo, Xianbin Liu, Ziqiang Zhao
Applications: Haomin Liu, Haoyu Ji, Donghui Shen
Art Designer: Yifu Wang, Zhewen Le
Project Lead: Guofeng Zhang, Xiaoyu Zhang
Acknowledgments
We thank Zhenzhou Fan, Jingfan Huang, Zhiwei Huang, Jingbo Liu, Chaoran Tian, Shao Zhao (listed alphabetically by last name) for their valuable discussions and assistance.

### 06 May 2026 v3

@@
InSpatio Team (alphabetical order) Donghui Shen, Guofeng Zhang, Haomin Liu, Haoyu Ji, Jialin Liu, Jing Guo, Nan Wang, Siji Pan, Weihong Pan, Weijian Xie, Xiaojun Xiang, Xiaoyu Zhang, Xianbin Liu, Yifu Wang, Yipeng Chen, Zhewen Le, Zhichao Ye, Ziqiang Zhao
@@ 3 Evaluation
To enable real-time interaction, we apply distribution matching distillation to obtain a distilled model for low-latency deployment. In its baseline configuration at 512×512 resolution, InSpatio-WorldFM achieves around 25 FPS on a single H-series GPU through engineering optimizations such as KV-cache management and efficient VAE latent caching. Moreover, due to its low GPU memory footprint, the model can run at 10 FPS on an RTX 4090. While distillation introduces a quality-speed trade-off, the perceptual difference remains minimal in practice. Visual comparison between the fundamental model and the distilled variant reveals that InSpatio-WorldFM successfully retains spatial coherence and geometric accuracy. The accelerated model continues to generate plausible content across diverse camera trajectories without noticeable degradation in structural consistency or the emergence of visual artifacts, confirming that our distillation approach effectively balances real-time performance with generation fidelity.